OpenStudy (anonymous):

(-1.5x-1.5y=-60) + (0.75x+1.5y=40) Solve for x

OpenStudy (uri):

Remove the paretheses. -1.5x-1.5y=-60+0.75x+1.5y=40

OpenStudy (anonymous):

y~~40.-x

OpenStudy (anonymous):

uri, would you show me how to solve this?

OpenStudy (uri):

@Jamierox4ev3r wants to